OpenStudy (anonymous):

Arlene is showing her work in simplifying −7.9 + 8.2 − 3.4 + 2.1. Identify any errors in her work or in her reasoning. Write feedback to Arlene explaining the mistake she made and how to correct her work. Step Work Reasoning −7.9 + 8.2 − 3.4 + 2.1 Original problem 1 −7.9 + 8.2 + (−3.4) + 2.1 Additive inverse 2 −7.9 + (−3.4) + 8.2 + 2.1 Associative property 3 (−7.9 + (−3.4)) + (8.2 + 2.1) Commutative property 4 −11.3 + 10.3 Simplify 5 −1 Simplify

OpenStudy (anonymous):

Keep doing each step manually, and see where it slips up: −7.9+8.2−3.4+2.1=−1 −7.9+8.2+(−3.4)+2.1=−1 −7.9+(−3.4)+8.2+2.1=−1 (−7.9+(−3.4))+(8.2+2.1)=−1 −11.3+10.3=−1

OpenStudy (anonymous):

So if you haven't noticed there are no errors

OpenStudy (anonymous):

ok but what about for Write feedback to Arlene explaining the mistake she made and how to correct her work.

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

Try answering the question and see where an error shows up, like @ChrystalC said :)

OpenStudy (anonymous):

she said there are no errors @RainbowBrony555

OpenStudy (anonymous):

@UnkleRhaukus can you help?

OpenStudy (unklerhaukus):

i don't see any errors

OpenStudy (anonymous):

Theres no errors, @KAKES1967

OpenStudy (anonymous):

ok so how would i go about this part though > Write feedback to Arlene explaining the mistake she made and how to correct her work. i would do nothing right?

OpenStudy (anonymous):

I think so

oregonduck (oregonduck):

Arlene is showing her work in simplifying −7.9 + 8.2 − 3.4 + 2.1. Identify any errors in her work or in her reasoning. Write feedback to Arlene explaining the mistake she made and how to correct her work. Step Work Reasoning −7.9 + 8.2 − 3.4 + 2.1 Original problem 1 −7.9 + 8.2 + (−3.4) + 2.1 Additive inverse 2 −7.9 + (−3.4) + 8.2 + 2.1 Associative property 3 (−7.9 + (−3.4)) + (8.2 + 2.1) Commutative property 4 −11.3 + 10.3 Simplify 5 −1 Simplify the properties in steps 2 & 3 are misidentified